In 2018, the West $10,000,000. gate Construction Company entered into a contract to construct a road for Santa Clara County for The road was completed in 2020. Information related to the contract is as follows: 2020 2018 2,542,000 $3,772,000 $2,074,600 5, 658,000 1,886,000 2,020,000 4,294,000 3,686,000 1,810,000 3,800,000 4,390,000 2019 Cost incurred during the year Estimated costs to complete as of year-end Billings during the year Cash collections during the year Westgate recognizes revenue over time according to percentage of completion. Required: 1. Calculate the amount of revenue and gross profit (oss) to be recognized in each of the three years. (Do not round Intermediate calculations. Loss amounts should be indicated with a minus sign.) 2018 2019 2020 Revenue Gross profit (loss)--
Answer #1

Solution

2018

2019

2020

Revenue

$       31,00,000.00

$ 46,00,000.00

$      23,00,000.00

Gross profit

$         5,58,000.00

$    8,28,000.00

$        2,25,400.00

Working

Working

2018

2019

2020

A

Contract Price

$    1,00,00,000

$        1,00,00,000

$       1,00,00,000

B

Cost Incurred to Date

$        25,42,000

$            63,14,000

$           83,88,600

C

Estimated cost yet to be incurred to complete the contract

$        56,58,000

$            18,86,000

$                          -  

D = B+C

Total Cost

$        82,00,000

$            82,00,000

$           83,88,600

E = (D/B) x 100

% of Complition

31%

77%

100%

F = A x E

Revenue to date

$ 31,00,000.00

$      77,00,000.00

$ 1,00,00,000.00

G

Revenue of Previous year

$                       -  

$      31,00,000.00

$     77,00,000.00

H = F - G

Net Revenue this year

$ 31,00,000.00

$      46,00,000.00

$     23,00,000.00

I

Cost to date

$ 25,42,000.00

$      63,14,000.00

$     83,88,600.00

J

Cost to date of previous year

$                       -  

$      25,42,000.00

$     63,14,000.00

K = I - J

Net Cost for the year

$ 25,42,000.00

$      37,72,000.00

$     20,74,600.00

L = H - K

Gross Profits

$    5,58,000.00

$        8,28,000.00

$       2,25,400.00